# Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation

### v4

@@ 4.4 Qualitative Analysis
As shown in  Fig. 6, compared to DUET, our SUSA agent is able to accurately stop near the “living room with a couch.” In contrast, the DUET agent, which relies exclusively on RGB environment representations, is prone to making elusive actions, hindering its ability to accurately follow instructions and reach the target location. This manifest that, thanks to the salient landmarks provided by the SUSA architecture, the agent is better able to navigate to the target location by grounding textual instructions.
@@ 5 Conclusion
We present the SUSA architecture, which exploits complementary semantic-spatial representations beyond RGB to facilitate environmental understanding and instruction grounding. Specifically, we propose the TSU module, enabling the agent to identify the most instruction-relevant semantic. The DSP module enhances spatial awareness by grounding depth exploration maps. We further employ contrastive learning to precisely align hybrid environmental representations with instructions. Experimental results on the R2R, REVERIE, and SOON benchmarks demonstrate the effectiveness and generalization of our approach. Future work will enrich informative environmental representations to advance VLN tasks in continuous environments.
Supplementary Material
6 Detailed Experimental Setups
Metrics. We categorize the metrics into three types: Navigation, Objects Grounding, and Instruction Following. Detailed calculations of these metrics can be found in  Tab. 7. For Navigation metrics, Trajectory Length (TL): Average path length in meters. It is worth noting that since the trajectory needs to match the ground truth trajectory described by the instruction, strictly speaking, the trajectory length is not necessarily the smaller, the better. Success Rate (SR): Proportion of paths where the agent reaches within 3 meters of the target location. Oracle Success Rate (OSR): The proportion of trajectories where at least one node is within 3 meters of a corresponding node in the reference trajectory. Navigation Error (NE): Average final distance (meters) between the agent and the target. Success Rate weighted by Path Length (SPL): Prioritizes success for shorter paths, which normalizes the success rate by trajectory length. For Instruction Following metrics, we apply Normalized Dynamic Time Warping (nDTW) and Success Rate Weighted by Dynamic Time Warping (sDTW). The former measures how well the VLN agent could follow the instruction, while the latter considers nDTW for success cases. For Objects Grounding metrics: Remote Grounding Success (RGS): Proportion of instructions where the agent correctly identifies the target object. RGS weighted by Path Length (RGSPL): Prioritizes correct object identification for shorter paths.
Metric $\uparrow$ $\downarrow$ Definition Dataset
@@ 6 Detailed Experimental Setups
nDTW $\uparrow$ $\exp\left(-\frac{\min_{W\in\mathcal{W}}\sum_{(i_{k},j_{k})\in W}d(p_{i_{k}},q_{j_{k}})}{|P|\cdot d_{th}}\right)$ R2R
sDTW $\uparrow$ $SR(P,G)*nDTW(P,G)$
@@
RGS $\uparrow$ $\mathbb{I}\Big{(}\max_{j}IoU(B_{i}^{p},B_{j}^{g})\geq 0.5\Big{)}$ REVERIE SOON
@@
Table 7: Common VLN metrics, categorized into three types: Navigation, Instruction Following, and Object Grounding. The symbols $\uparrow$ and $\downarrow$ signify if a higher or lower metric value is preferable, respectively. P represents the paths taken by the agent, while G stands for the ground truth trajectories. The function $\mathbb{I}(\cdot)$ serves as an indicator, with $d(n,P)$ measuring the distance from node $n$ to path $P$ , and $d_{th}$ indicating the distance threshold. $W=w_{1..|W|}$ is a warping with $w_{k}=(i_{k},j_{k})\in[1:|P|],[1:|G|]$, respecting the step-size.$B^{p}_{i}$ means the i-th object’s bounding box in agent path (or trajectory).
Pre-training Tasks. Following prior works [6, 5, 30], during pre-training, we employ a suite of multimodal pre-training tasks, including MLM (Masked Language Modeling), MRC (Masked Region Classification), SAP (Single Action Prediction), and OG (Object Grounding). OG is specifically designed for object grounding using the REVERIE and SOON datasets. MLM and MRC aim to establish a robust vision-language alignment. SAP enables the agent to learn preliminary navigation knowledge by predicting the next action during pre-training. However, as navigation is inherently a sequential decision-making process, SAP can only learn static, single-step decisions.
Dataset Train Validation Seen Validation Unseen Test Unseen VLN task
@@
R2R [2] 14039 61 1021 56 2349 11 4173 18 Fine-grained
REVERIE [41] 10466 60 1423 46 3521 10 6292 16 Goal-oriented
SOON [64] 2779 34 113 2 339 5 615 14 Goal-oriented
@@
Dataset Details. We evaluate the navigation capabilities of the agent on three distinct VLN datasets: the fine-grained R2R [2], as well as the goal-oriented REVERIE [41] and SOON [64] datasets, as shown in Tab. 8. Each dataset is divided into train, validation seen, validation unseen, and test unseen splits. REVERIE contains approximately 20k high-level instructions describing target locations and objects, with an average instruction length of 21 words. Given predefined object bounding boxes within each panorama, the agent identifies the correct bounding box upon reaching the navigation goal. Unlike the R2R dataset, in the REVERIE task, the agent must identify the correct bounding box from the candidates once the navigation target is reached. R2R includes 22k step-by-step instructions, with an average instruction length of 32 words. The R2R dataset focuses solely on the navigation task, providing detailed path planning for the agent in the instructions. SOON is another goal-oriented dataset, containing about 5k instructions, each averaging 47 words in length.
We evaluated the experimental results of the unseen test split on the public online leaderboards [^1] https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate. We assure the readers that we will make our code and model checkpoints publicly available.
7 Additional Experiments
Validation Unseen
Methods TL SR SPL RGS RGSPL
BEVBert [1] - 51.78 36.37 34.71 24.44
GridMM [56] 23.20 51.37 36.47 34.57 24.56
SEAT[53] - 49.45 35,51 32.83 23.14
SUSA(DP) 24.43 52.54 37.59 35.47 25.27
Depth SUSA(DEM) 22.88 52.00 39.21 35.05 26.54
KESU [15] 22.93 50.36 35.15 35.02 24.17
KERM [31] 21.85 50.44 35.38 34.51 24.45
AACL [33] 22.11 49.42 33.54 33.31 22.49
CONSOLE [34] - 50.07 34.40 34.05 23.33
Text SUSA(TP) 23.40 53.17 38.31 37.18 26.73
@@ 7 Additional Experiments
7.1 Comparison with Previous Approaches
We demonstrate the effectiveness of the proposed method through comparisons with related approaches. As shown in  Tab. 9, experiments reveal that directly grounding depth information—whether using depth panoramas (DP) or depth exploration maps (DEM) as input—outperforms prior depth-based methods [1, 56, 53]. Notably, using depth exploration maps leads to better navigation performance than depth panoramas. Similarly, directly aligning textual semantics with instructions via textual panoramas (TP) surpasses methods [15, 31, 34] that integrate textual semantics into environmental representations. This underscores the superiority of independently grounding each modality, improving the interpretability of each modality’s contribution.
Figure 7: The weights for hybrid representations in the HRA module change during training (see Eq. 6).
@@ 7.1 Comparison with Previous Approaches
7.2 Fusion Weights of Hybrid Representations.
We quantitatively illustrate the weights (see Eq. 6) assigned to each branch of the HRA module during training (in Fig. 7) and inference (in Tab. 10). These weights are normalized between 0 and 1, with higher values denoting stronger dependency on the corresponding environmental representation. As shown in Fig. 7, initially, all four representations have approximately equal weights of 0.25 during training, reflecting random initialization and balanced contributions. Over the course of training, the global RGB representation weight $\beta_{4}$ increases, while those of the textual semantic $\beta_{1}$ and depth $\beta_{3}$ modalities decrease. This trend arises because textual semantics and depth provide sparse yet precise information, which is particularly beneficial in the early stages of training. Despite the increasing dominance of the RGB modality, the agent’s reliance on textual semantics and depth remains above 20%, underscoring their essential role in navigation. Tab. 10 also demonstrates that the agent effectively utilizes the introduced depth spatial and textual semantic information during inference, even though it exhibits a stronger reliance on the RGB modality.
@@
Semantic $\beta_{1}$ RGB $\beta_{2}$ Depth $\beta_{3}$ RGB $\beta_{4}$
Validation Seen 0.2147 0.2458 0.2239 0.3156
Validation Unseen 0.2151 0.2662 0.2216 0.2971
Table 10: Variation in fusion weights (in Eq. 6) of hybrid representations during inference on R2R. $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HRA module, respectively.
7.3 Quality of Generated Salient Landmarks
The quality of the extracted salient landmarks is crucial for the agent to accurately understand the textual semantics of the environment. In Fig. 8, we compute the similarity between the top-50 most frequently occurring landmarks in our generated salient landmarks and those from R2R [2] and REVERIE [41] to evaluate the quality of the generated salient landmarks. While the types and frequencies of these generated landmarks is different from those in the datasets, they can contain and convey most similar sematic (similarities surpass 0.5). Notably, the plug-and-play CLIP text encoder excels in understanding image-related descriptions, while BERT is more adept at comprehending natural language sentences. Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks.
Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HRA module on R2R validation unseen split.
7.4 Analysis of the Contrastive Loss Weight
@@
[…]
1 13.41 3.30 79.86 71.43 60.63
2 13.85 3.08 80.42 72.80 62.71
3 13.78 3.06 81.65 72.92 63.16
4 12.18 3.06 79.99 73.05 64.85
5 14.58 3.08 82.08 73.95 62.84
7.5 Overall Design on R2R
We further supplemented the performance of each key component of SUSA on the R2R task. As shown in Tab. 6, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL). The TSU module is better at understanding textual semantics, resulting in higher success rates. After integrating the DSP and TSU modules, the agent’s performance improved significantly (e.g., SR=73.05, SPL=64.85). When the HRA module is further introduced, the navigation success rate on R2R increases to 73.95, indicating that contrastive learning further enhances instruction-environment alignment. However, its SPL metric remains suboptimal, suggesting that the contrastive learning strategy in the HRA module may excel at accurately aligning the environment with instructions but lacks the capability to optimize the navigation path length.
7.6 Computing Efficiency vs. Navigation Performance
@@ 7.5 Overall Design on R2R
As Fig. 10 illustrated, we compare the Giga Floating-point Operations (GFLOPs) and navigation performance (i.e., SR/SPL, which are more important metrics) of different VLN models, including DUET [6] , DSRG [50], HAMT [5] and our SUSA. To ensure a fair comparison across all methods, we performed single-step forward inference with a batch size of 8, an instruction length of 44, and 6 exploration map nodes. Compared to the baseline DUET, our approach introduces spatial and semantic environmental information, which inevitably increasing computational cost. However, benefiting from the partial pretraining strategy and a shared panorama encoder for both the depth and RGB environmental information, our SUSA architecture notably enhances agent navigation performance with modest computational resources. The frozen BLIP-2 and CLIP models provide the agent with off-the-shelf textual semantics and are excluded from the training or inference.
8 Qualitative Examples
Failure cases. We provide an in-depth examination of failure cases, aiming to offer valuable insights for future model improvements. Fig. 11 displays the difference between the navigation trajectory predicted by the proposed SUSA and the ground truth trajectory. We found that, although textual semantics can provide the agent with a richer understanding, it may introduces ambiguity in certain scenarios. Concretely, when two identical salient landmarks appear in different scenes (e.g., the staircase and bathroom in Fig. 11), the agent may struggle to distinguish the correct navigable node, leading to erroneous navigation decisions.
@@ 8 Qualitative Examples
Figure 11: Top-down trajectories visualization of failed cases. The corresponding instructions are shown at the bottom. Partial views of final stopping position and corresponding landmark captions of ground truth and SUSA are also provided
9 Limitations and Future Work

### v5

@@ 4.4 Qualitative Analysis
As shown in Fig. 6, compared to DUET, our SUSA agent is able to accurately stop near the “living room with a couch.” In contrast, the DUET agent, which relies exclusively on RGB environment representations, is prone to making elusive actions, hindering its ability to accurately follow instructions and reach the target location. This manifests that, thanks to the salient landmarks provided by the SUSA architecture, the agent is better able to navigate to the target location by grounding textual instructions.
@@ 5 Conclusion
We present the SUSA architecture, which hierarchical exploits semantic-spatial representations beyond RGB to facilitate environmental understanding and instruction grounding. Specifically, we propose the TSU module, enabling the agent to identify the most instruction-relevant textual semantic. The DSP module complementarily enhances spatial awareness by incrementally constructing and grounding depth exploration maps. Experimental results on the R2R, REVERIE, SOON and R2R-CE benchmarks demonstrate the navigational superiority of our approach. Benefiting from hierarchical representation enrichment and alignment, SUSA exhibits excellent generalization performance on unseen environments and modality interpretability.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants Nos. 62202139, 62573292 and 62172138, and also supported by the Fundamental Research Funds for the Central Universities under Grants Nos.JZ2025HGTB0226, JZ2024HGTG0310 and Young Elite Scientist Sponsorship Program (No. YESS20220409). The computations were performed on the High-Performance Computing (HPC) Platform of Hefei University of Technology.
Appendices
This appendix provides additional materials to complement the main paper. Specifically:
  • Section A presents detailed descriptions of metrics, dataset details and pre-training tasks.
  • Section B offers extended experimental results, including: (1) comparisons with related approaches, (2) analysis of the learned fusion weights across modalities, (3) evaluation of the quality of generated salient landmarks, (4) ablations on the contrastive loss weight, (5) an overview of the overall design on the R2R benchmark, and (6) a discussion on the trade-off between computational efficiency and navigation performance.
  • Section C showcases additional qualitative examples to analyse the agent’s behavior.
  • Section D discusses the limitations of our current approach and outlines potential directions for future work.
Appendix A Detailed Experimental Setups
Metrics. We categorize the metrics into three types: Navigation, Objects Grounding, and Instruction Following. Detailed calculations of these metrics can be found in Table 7. For Navigation metrics, Trajectory Length (TL): Average path length in meters. It is worth noting that since the trajectory needs to match the ground truth trajectory described by the instruction, strictly speaking, the trajectory length is not necessarily the smaller, the better. Success Rate (SR): Proportion of paths where the agent reaches within 3 meters of the target location. Oracle Success Rate (OSR): The proportion of trajectories where at least one node is within 3 meters of a corresponding node in the reference trajectory. Navigation Error (NE): Average final distance (meters) between the agent and the target. Success Rate weighted by Path Length (SPL): Prioritizes success for shorter paths, which normalizes the success rate by trajectory length. For Instruction Following metrics, we apply Normalized Dynamic Time Warping (nDTW) and Success Rate Weighted by Dynamic Time Warping (sDTW). The former measures how well the VLN agent could follow the instruction, while the latter considers nDTW for success cases. For Objects Grounding metrics: Remote Grounding Success (RGS): Proportion of instructions where the agent correctly identifies the target object. RGS weighted by Path Length (RGSPL): Prioritizes correct object identification for shorter paths.
Metric $\uparrow$ $\downarrow$ Definition
@@ Appendix A Detailed Experimental Setups
nDTW $\uparrow$ $\exp\left(-\frac{\min_{W\in\mathcal{W}}\sum_{(i_{k},j_{k})\in W}d(p_{i_{k}},q_{j_{k}})}{|P|\cdot d_{th}}\right)$
sDTW $\uparrow$ SR(P,G)nDTW(P,G)SR(P,G)*nDTW(P,G)
@@
RGS $\uparrow$ $\mathbb{I}\Big(\max_{j}IoU(B_{i}^{p},B_{j}^{g})\geq 0.5\Big)$
@@
[…]
Dataset Train Val Seen Val Unseen Test Unseen VLN task
@@
R2R (Anderson et al. 2018) 14039 61 1021 56 2349 11 4173 18 Fine-grained
R2R-CE (Krantz et al. 2020) 10819 61 778 53 1839 11 3408 18 Fine-grained
REVERIE (Qi et al. 2020) 10466 60 1423 46 3521 10 6292 16 Goal-oriented
SOON (Zhu et al. 2021) 2779 34 113 2 339 5 615 14 Goal-oriented
@@
Dataset Details. We primarily evaluate the agent’s navigation capabilities on three distinct VLN datasets: the fine-grained R2R (Anderson et al. 2018), and the goal-oriented REVERIE (Qi et al. 2020) and SOON (Zhu et al. 2021). To further assess generalization in continuous environments (Savva et al. 2019), we extend our evaluation to the R2R-CE (Krantz et al. 2020) task, as shown in Table 8. Each dataset is divided into train, validation seen, validation unseen, and test unseen splits. REVERIE contains approximately 20k high-level instructions describing target locations and objects, with an average instruction length of 21 words. Given predefined object bounding boxes within each panorama, the agent identifies the correct bounding box upon reaching the navigation goal. Unlike the R2R dataset, in the REVERIE task, the agent must identify the correct bounding box from the candidates once the navigation target is reached. R2R includes 22k step-by-step instructions, with an average instruction length of 32 words. The R2R dataset focuses solely on the navigation task, providing detailed path planning for the agent in the instructions. In R2R-CE, the average distance between navigable nodes is significantly shorter compared to R2R (0.19 vs. 2.25 meters), while the number of navigation steps is substantially higher (55 vs. 4–6). SOON is another goal-oriented dataset, containing about 5k instructions, each averaging 47 words in length.
Pre-training Tasks. Following prior works (Chen et al. 2022b, 2021), during pre-training, we employ a suite of multimodal pre-training tasks, including MLM (Masked Language Modeling), MRC (Masked Region Classification), SAP (Single Action Prediction), and OG (Object Grounding). OG is specifically designed for object grounding using the REVERIE and SOON datasets. MLM and MRC aim to establish a robust vision-language alignment. SAP enables the agent to learn preliminary navigation knowledge by predicting the next action during pre-training. However, as navigation is inherently a sequential decision-making process, SAP can only learn static, single-step decisions. When transferring to the R2R-CE task, the agent—serving as part of the planner module—is not trained from scratch in continuous environments, but only fine-tuned.
We evaluated the experimental results of the unseen test split on the public online leaderboards [^1] https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate.. We assure the readers that we will make our code and model checkpoints publicly available.
Appendix B Additional Experiments
Val Unseen
Methods SR SPL RGS RGSPL
BEVBert 51.78 36.37 34.71 24.44
GridMM 51.37 36.47 34.57 24.56
SEAT 49.45 35,51 32.83 23.14
SUSA(DP) 52.54 37.59 35.47 25.27
Depth SUSA(DEM) 52.00 39.21 35.05 26.54
KESU 50.36 35.15 35.02 24.17
KERM 50.44 35.38 34.51 24.45
AACL 49.42 33.54 33.31 22.49
CONSOLE 50.07 34.40 34.05 23.33
Text SUSA(TP) 53.17 38.31 37.18 26.73
@@ Appendix B Additional Experiments
B.1 Comparison with Related Approaches
We demonstrate the effectiveness of the proposed method through fair comparisons with related approaches. As shown in Table 9, experiments reveal that directly grounding depth information—whether using depth panoramas (DP) or depth exploration maps (DEM) as input—outperforms prior depth-based methods (An et al. 2023; Wang et al. 2023c, 2024b). Notably, using depth exploration maps leads to better navigation performance than depth panoramas. Similarly, directly aligning textual semantics with instructions via textual panoramas (TP) surpasses methods (Gao et al. 2024; Li et al. 2023b; Lin et al. 2024) that integrate textual semantics into environmental representations. This underscores the superiority of independently grounding each modality, improving the interpretability of each modality’s contribution.
Figure 7: The weights for hybrid representations in the HAP module change during training (see Eq.  6).
@@ B.1 Comparison with Related Approaches
B.2 Fusion Weights of Hybrid Representations.
We quantitatively illustrate the weights (see Eq. 6) assigned to each branch of the HAP module during training (in Fig. 7) and inference (in Table 10). These weights are normalized between 0 and 1, with higher values denoting stronger dependency on the corresponding environmental representation. As shown in Fig. 7, initially, all four representations have approximately equal weights of 0.25 during training, reflecting random initialization and balanced contributions. Over the course of training, the global RGB representation weight $\beta_{4}$ increases, while those of the textual semantic $\beta_{1}$ and depth $\beta_{3}$ modalities decrease. This trend arises because textual semantics and depth provide sparse yet precise information, which is particularly beneficial in the early stages of training. Despite the increasing dominance of the RGB modality, the agent’s reliance on textual semantics and depth remains above 20%, underscoring their essential role in navigation. Table 10 also demonstrates that the agent effectively utilizes the introduced depth spatial and textual semantic information during inference, even though it exhibits a stronger reliance on the RGB modality.
@@
$\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{4}$
Val Seen 0.2147 0.2458 0.2239 0.3156
Val Unseen 0.2151 0.2662 0.2216 0.2971
Table 10: Variation in fusion weights (in Eq. 6) of hybrid representations during inference on R2R. $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HAP module, respectively.
Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HAP module on R2R validation unseen split.
B.3 Quality of Generated Salient Landmarks
The quality of the extracted salient landmarks is crucial for the agent to accurately understand the textual semantics of the environment. In Fig. 8, we compute the similarity between the top-50 most frequently occurring landmarks in our generated salient landmarks and those from R2R (Anderson et al. 2018) and REVERIE (Qi et al. 2020) to evaluate the quality of the generated salient landmarks. While the types and frequencies of these generated landmarks is different from those in the datasets, they can contain and convey most similar sematic (similarities surpass 0.5). Notably, the plug-and-play CLIP text encoder excels in understanding image-related descriptions, while BERT is more adept at comprehending natural language sentences. Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks.
[…]
1 13.41 79.86 71.43 60.63
2 13.85 80.42 72.80 62.71
3 13.78 81.65 72.92 63.16
4 12.18 79.99 73.05 64.85
5 14.58 82.08 73.95 62.84
B.4 Analysis of the Contrastive Loss Weight
@@
B.5 Overall Design on R2R
We further supplemented the performance of each key component of SUSA on the R2R task. As shown in Table 11, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL). The TSU module is better at understanding textual semantics, resulting in higher success rates. After integrating the DSP and TSU modules, the agent’s performance improved significantly (e.g., SR=73.05, SPL=64.85). When the HAP module is further introduced, the navigation success rate on R2R increases to 73.95, indicating that contrastive learning further enhances instruction-environment alignment. However, its SPL metric remains suboptimal, suggesting that the contrastive learning strategy in the HAP module may excel at accurately aligning the environment with instructions but lacks the capability to optimize the navigation path length.
B.6 Computing Efficiency vs. Navigation Performance
@@ B.5 Overall Design on R2R
As 10 illustrated, we compare the Giga Floating-point Operations (GFLOPs) and navigation performance (i.e, SR/SPL, which are more important metrics) of different VLN models, including DUET (Chen et al. 2022b) , DSRG (Wang et al. 2023a), HAMT (Chen et al. 2021) and our SUSA. To ensure a fair comparison across all methods, we performed single-step forward inference with a batch size of 8, an instruction length of 44, and 6 exploration map nodes. Compared to the baseline DUET, our approach introduces spatial and semantic environmental information, which inevitably increasing computational cost. However, benefiting from the partial pretraining strategy and a shared panorama encoder for both the depth and RGB environmental information, our SUSA architecture notably enhances agent navigation performance with modest computational resources. The frozen BLIP-2 and CLIP models provide the SUSA agent with off-the-shelf textual semantics and are excluded from the training or inference.
Appendix C Qualitative Examples
Failure cases. We provide an in-depth examination of failure cases, aiming to offer valuable insights for future model improvements. 11 displays the difference between the navigation trajectory predicted by the proposed SUSA and the ground truth trajectory. We found that, although textual semantics can provide the agent with a richer understanding, it may introduces ambiguity in certain scenarios. Concretely, when two identical salient landmarks appear in different scenes (e.g., the staircase and bathroom in 11), the agent may struggle to distinguish the correct navigable node, leading to erroneous navigation decisions.
@@ Appendix C Qualitative Examples
Figure 11: Top-down trajectories visualization of failed cases. Partial views of final stop position and corresponding landmark captions are also provided.
Appendix D Limitations and Future Work